OpenStudy (anonymous):

Simplify product using same rules for exponets (9^7) ^3

OpenStudy (zale101):

multiply the exponents

OpenStudy (anonymous):

the rule this time is: \[a^{n^m}=a^{mn}\]

OpenStudy (anonymous):

a=9 m=7 n=3

OpenStudy (anonymous):

So it would be 189?

OpenStudy (anonymous):

I am slow at math lol

OpenStudy (anonymous):

it would be \[9^{7*3}=9^{21}\]

OpenStudy (anonymous):

Thank you what is the simplified expression for (6^2)^7

OpenStudy (anonymous):

you give me the answer and I tell you if it's right or no ? use the same rule replace a,m and n just try

OpenStudy (anonymous):

Okay is it 6^ 5

OpenStudy (anonymous):

nop , what 7x2 is ?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

6^14

OpenStudy (anonymous):

so it's \[6^{2*7}=6^{14}\]

OpenStudy (anonymous):

you see ?

OpenStudy (anonymous):

yes :) THANKS

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

\[\left(9^7\right)^3=\left(\left(3^2\right)^7\right)^3=3^{42} \]

OpenStudy (anonymous):

\[\sqrt[42]{109418989131512359209}=3 \] from Mathematica which is capable of unlimited precision arithmetic.